what Celsius temperature is needed to change the volume of a gas from 1500 ml at 25c to 2100ml
Answer #1

We know that PV = nRT;

we can write for two as P1V1 = nRT1 & P2V2 = nRT2 at constant pressure these two combined equations becomes Volume in Liters and Temperature in Kelvin only

V1/V2 = T1/T2   =>

T1 = (V1/V2) X T2

     = (2.1/1.5) X 298

     = 417.2 K

Temperature needed = 144.2 oC